X=(31 over 6)^3/2 * $1,000,000
Answer by checkley77(12844)

X=(31 over 6)^3/2 * $1,000,000
Do you mean [(31/6)^3/2]1,000,000 ?
31/6=5.16667
5.16667^3/2=11.744
11.744*1,000,000=11,744,000